2. The Tl in a 9.76-g sample of rodenticide was oxidized to the trivalent state and treated with an unmeasured excess of Mg/EDTA solution. The reaction is Tl3+ + MgY2- → TlY- + Mg2+
.Titration of the liberated Mg2+ required 13.34 mL of 0.03560 M EDTA. Calculate the % of Tl2SO4 (504.8 g/mol) in the sample.
1
Expert's answer
2020-06-04T10:36:51-0400

Solution:

H2Y2- = EDTA

The reactions are:

(1) Tl3+ + MgY2- → TlY- + Mg2+ 

(2) Mg2+ + H2Y2- → MgY2- + 2H+


According to the (2) equation: n(EDTA) = n(Mg2+)

Each mole of EDTA reacts with one mole of Mg2+.

n(EDTA) = Molarity of EDTA × Volume of solution

n(EDTA) = (0.03560  M) × (0.01334 L) = 0.0004749 mol

n(Mg2+) = n(EDTA) = 0.0004749 mol


According to the (1) equation: n(Mg2+) = n(Tl3+)

For every one mole of Mg2+, there one mole of Tl3+

n(Tl3+) = n(Mg2+) = 0.0004749 mol


(3) Tl2SO4 → 2Tl3+

According to the (3) equation: n(Tl2SO4) = n(Tl3+)/2

There are two thallium ions per one molecule of Tl2SO4.

n(Tl2SO4) = n(Tl3+)/2 = (0.0004749 mol) / 2 = 0.00023745 mol


Moles of Tl2SO4 = n(Tl2SO4) = Mass of Tl2SO4 / Molar mass of Tl2SO4

The molar mass of Tl2SO4 is 504.8 g/mol.

So, the mass of thallium sulphate is:

Mass of Tl2SO4 = n(Tl2SO4) × M(Tl2SO4)

Mass of Tl2SO4 = (0.00023745 mol) × (504.8 g/mol) = 0.11986 g


Finally, the percentage of thallium sulphate (Tl2SO4) in the sample is:

% of Tl2SO4 = m(Tl2SO4) / m(sample) = (0.11986 g) / (9.76 g) = 0.01228 or 1.23%

% of Tl2SO4 = 1.23 %


Answer: 1.23% of Tl2SO4 in the sample.